### v2

@@ Abstract
The common observation of a $1/f^{\alpha}$’ power spectrum with $\alpha<2$ constitutes one of the enduring mysteries of condensed matter physics. Here it is shown that a $1/f^{\alpha}$ power spectrum, with $\alpha=3/2$, can arise when an ensemble of two–level systems is coupled to a heat bath by means of a system of Bosonic quasiparticles. The model considered is the classic model of Faughnan and Strandberg of the phonon bottleneck, and the anomalous relaxation is associated with an approximate non-equilibrium steady state of the phonons maintained by slow spin relaxation. It is shown that a frequency-dependent susceptibility can be defined in the steady state and that the prediction $\alpha=3/2$, or the equivalent stretched exponential relaxation with exponent $\beta=2-\alpha=1/2$, is consistent with existing experimental data. These results give an insight into the origins of anomalous relaxation in condensed matter and the practical behaviour of two-level systems.
@@
Exponential, or Debye, relaxation has a power spectrum that tends to $1/f^{2}$ at high frequency $f$. However many real condensed matter systems – conducting [1, 2], dielectric [3, 4, 5] and magnetic [6, 7, 8, 9, 10, 11] – show power law behavior $\sim 1/f^{\alpha}$ with an exponent $\alpha$ that lies in the range $1\leq\alpha<2$. While a number of particular mechanisms for this behavior have been established (see e.g. Refs. [6, 7, 8, 9, 10] for the magnetic case), in general it remains a mystery as to why $1/f^{\alpha}$ behavior is so widespread. A common explanation [1] is that the power law is constructed from a wide range of exponential timescales, but in physical terms the range of timescales required (many decades) is usually unrealistic. An exception might be where the time scales are associated with an underlying broad range of length scales, as occurs in a critical or fractal system, but such instances cannot account for the majority of observations. This leads one to suspect that there are more fundamental causes at play, arising in non-equilibrium physics and quantum dynamics. In this context, simple and realistic models and experiments that produce well-defined exponents $\alpha$ are of evident interest to signpost the way to fundamental causes.
This paper reconsiders the canonical model of the ‘phonon bottleneck’ due to Faughnan and Strandberg [12]. The model exemplifies a simple non-equilibrium, and essentially quantum mechanical, model that describes an ensemble of two-level systems (spins) connected to a heat bath by an intervening system of Bosonic quasiparticles (phonons). In a certain parameter regime the model describes a non-equilibrium steady state, and it is shown here that the steady state is associated with almost perfect $1/f^{3/2}$ power spectrum. This result, derived below, is illustrated in Fig. 1 where the exact solution for $|\chi^{\prime\prime}|/\omega$ (the ratio of the imaginary part of the magnetic susceptibility to the angular frequency, a measure of the power spectrum [13]) is compared to the approximation :
$\frac{|\chi^{\prime\prime}|}{\omega}=\frac{\chi_{T}}{\sigma^{-1}+\sqrt{2\sigma}\,\omega^{3/2}}$ (1)
(where $\chi_{T}$ is the isothermal susceptibility), and a $1/f^{3/2}$ power spectrum (given $f=\omega/2\pi$) is observed over many decades of frequency.
Figure 1: Exact power spectrum (blue line) compared with Eq. 1 (magenta dashed line). Here $\sigma=10^{5}$ and a power law region with $\alpha=3/2$ extends over $1/\sigma<\omega<\sigma$. The two curves are almost coincident on a log–log scale, illustrating a rare instance where a $1/f^{\alpha}$’ power spectrum can be calculated analytically (here $\alpha=3/2$ and $f=\omega/2\pi$).
The model considers a dilute system of two-state spins that relax by directly emitting or absorbing phonons in a narrow band around the resonant frequency. Owing to ‘quantum freezing’ of phonon modes at low temperatures, the phonon specific heat becomes too small to maintain thermal equilibrium, hence the ‘bottleneck’. The phonons are treated quantum mechanically [14], leading to two coupled equations of motion.
@@
Here the variable $v$ is defined as
$v=\frac{(N_{i}^{0}-N_{j}^{0})-(N_{i}-N_{j})}{(N_{i}^{0}-N_{j}^{0})}$ (4)
where $N_{i}$ is the number of spins in the excited state, $N_{j}$ is the number of spins in the ground state, the superscripts represent equilibrium values and $N_{i}+N_{j}=N_{i}^{0}+N_{j}^{0}=N$, the number of spins. If $i$ denotes spin up, and $j$ denotes spin down, then the magnetic moment $\mathcal{M}=N_{i}-N_{j}$ and the equilibrium magnetic moment is $\mathcal{M}^{0}=N_{i}^{0}-N_{j}^{0}$. It follows that $v=(\mathcal{M}^{0}-\mathcal{M})/\mathcal{M}^{0},$ a normalised measure of the deviation of magnetic moment, or magnetization, from its equilibrium value. For example, if a field is suddenly applied to an equlibrium unmagnetized system, then $v$ will evolve from $v=1$ to $v=0$ as the new equilibrium is established at $\mathcal{M}=\mathcal{M}^{0}$.
The variable $u$ is defined as
@@
where the $p$’s are mean phonon occupation numbers at the frequency corresponding to the level splitting. Finally the two parameters are defined as follows. The parameter $a=\tau_{\rm s}/\tau_{\rm p}$ where $\tau_{\rm s}$ is a spin relaxation time and $\tau_{p}$ is a phonon relaxation time. The parameter $\sigma$ is defined by
@@
where the $c$’s are specific heats. A large value of $\sigma$ indicates a relatively small phonon specific heat: hence it is a measure of the strength of the ‘phonon bottleneck’. Note that, in this treatment, $\sigma$ and $a$ are dimensionless, so the time $t$ and angular frequency $\omega$ are also dimensionless, defined with respect to some reference time scale.
The situation of interest arises where $\sigma,a\gg 1$. The phonons are strongly bottlenecked but the spins relax relatively slowly, continually maintaining a far from equilibrium phonon population. This leads to an approximate steady state where the rate of change of the phonon population can be neglected. Thus $|du/dt|\ll a$, and the right hand side of Eq. 3 may be set to zero. This reduces Eqs. 2, 3 to a single equation depending only on $\sigma$:
@@
At short times $v\rightarrow 1$, but for sufficiently large $\sigma$, there appears to be a region at $1-v\gg 1/\sigma$ where this equation approximates:
@@
Correspondingly, the Fourier transform of this function at high frequency should approximate
$v(\omega)\approx\frac{1}{\sqrt{2\sigma}\omega^{3/2}},$ (9)
suggesting an equation analagous to Eq. 1 for $v(\omega)$.
These ideas are born out by numerical solution of the coupled equations. Using the example of $a=10^{2},\sigma=10^{5}$, Fig. 2 (upper) illustrates the solution for $v(t)$ compared to $1-\sqrt{2t/\sigma}$ and to $e^{-\sqrt{2t/\sigma}}$, which are very close approximations at small times, and the (exact) inverse Fourier transform [15] of the right hand side of Eq. 1, which captures the whole curve. Fig. 2 (lower) is the corresponding numerical Fourier transform on a log-log scale, confirming a slope very close to $3/2$. Fig. 3 illustrates the corresponding $u(t)$ compared to $\frac{d}{dt}e^{-\sqrt{2t/\sigma}}$ (see Eqs. 7 and 8) which is a close approximation at short time after transient effects have died out.
Figure 2: Relaxation of the magnetization difference: numerical solution to Eqs. 2, 3, for $a=10^{2},\sigma=10^{5}$. (a) the function $v(t)$ (blue line) compared to $1-\sqrt{2t/\sigma}$ (gray dotted line), $e^{-\sqrt{2t/\sigma}}$ (gray dashed line) and to the (exact) inverse Fourier transform of the right hand side of Eq. 1 (magenta dashed line). (b) Fast Fourier Transform of $v(t)$ (blue points) compared to a line of slope $3/2$ (orange line) on the log-log graph.
Figure 3: Relaxation of the phonon population: $u(t)$ numerically calculated in the example of Fig.2 (red dashed line) compared to the prediction $w\approx\frac{d}{dt}e^{-\sqrt{2t/\sigma}}$ (see Eq. 7). There is a close agreement once transient effects have died out at very short time.
Exact solutions to Eqs. 2, 3, 7, may be explored by rewriting the equations as a function of the magnetization $M=\mathcal{M}/V$, where $V$ is the volume. Eqs. 2,3 become
$\frac{dM}{dt}=(M^{0}-M)-\sigma uM$ (10)
$\frac{du}{dt}=a\left(1-\frac{M}{M^{0}}-u-\sigma u\frac{M}{M^{0}}\right)$ (11)
and Eq. 7 becomes:
$M^{0}-M=\frac{dM}{dt}\left(1+\frac{\sigma M}{M^{0}}\right).$ (12)
In the limit that Eq. 12 applies, now consider that the position of equilibrium is set by an applied field $H=H_{1}+he^{i\omega t}$. We have $M_{0}=\chi_{T}H$ where $\chi_{T}$ is the equilibrium isothermal susceptibility, and writing $M=M_{1}+m(\omega)e^{i\omega t}$ we may define a complex susceptibility $\chi(\omega)=\partial m/\partial h$. After differentiating Eq. 12 with respect to $h$, Eq. 12 becomes:
$\chi_{T}-\chi=i\omega\chi\left(1+\sigma\frac{\chi}{\chi_{T}}\right),$ (13)
from which the real and imaginary parts of $\chi$ can be separated:
$\chi^{\prime}=\frac{-\omega\chi_{T}+\chi_{T}(1+(2+16\sigma(1+\sigma))\omega^{2}+\omega^{4})^{1/4}\sin\phi}{2\sigma w}$ (14)
$\chi^{\prime\prime}=\frac{\chi_{T}-\chi_{T}(1+(2+16\sigma(1+\sigma))\omega^{2}+\omega^{4})^{1/4}\cos\phi}{2\sigma w}$ (15)
According to Eq. 15, for large $\sigma$, the function $|\chi^{\prime\prime}|/\omega$ tends to $\chi_{T}\sigma$ as $\omega\rightarrow 0$ and tends to
$\frac{|\chi^{\prime\prime}|}{\omega}\rightarrow\frac{\chi_{T}}{\sqrt{2\sigma}\,\omega^{3/2}},$ (16)
for large $\omega<\sigma$ (cf Eq. 1, Eq. 9). This justifies the close approximation of the exact form by Eq. 1, as displayed in Fig. 1. However, the high order terms in $\omega$ in Eq. 15 ensure that the true asymptotic behavior of $|\chi^{\prime\prime}|/\omega$ as $\omega\rightarrow\infty$ is $1/\omega^{2}$, with a crossover to this form at $\omega\sim\sigma$. Thus the anomalous $\alpha=3/2$’ scaling applies to the range $1/\sigma<\omega<\sigma$ and only the divergence of $\sigma$ would extend its range to infinity. Nevertheless, as illustrated in Fig. 4, for finite $\sigma$, the crossover is particularly sharp and has negligible effect on the power law scaling at smaller $\omega$. Were such behavior encountered in experiment, it would surely be interpreted as an asymptotic $1/\omega^{3/2}$ power law as one would have to access unfeasibly small values of $|\chi^{\prime\prime}|/\omega$ to see the crossover.
A similar analysis of the coupled equations shows that the existence of a frequency-dependent susceptibility $\chi(\omega)$ strictly implies either that $dw/dt=0$ (the case already discussed) or $a=0$, where Debye relaxation is recovered. Thus the two equations become:
$u=\frac{\chi_{T}-\chi-i\omega\chi}{\sigma\chi}$ (17)
$\frac{du}{dt}+u\left(1+\sigma\frac{\chi}{\chi_{T}}\right)=a\left(1-\frac{\chi}{\chi_{T}}\right)$ (18)
and a solution of the lower equation with $u=0$ at $t=0$ gives:
$u=\frac{\chi_{T}-\chi}{\chi_{T}+\sigma\chi}\left(1-e^{-at\left(1+\sigma\chi/\chi_{T}\right)}\right)$ (19)
which is consistent with Eq 17 (whose right hand side does not involve time) only in the limit $at\rightarrow\infty$. However, at $\omega\sim 1/t\ll a$, corrections to the limiting solution become exponentially small and the analytic form Eqs. 1, 15 should represent a close approximation to the solution of the coupled equations, as already observed (Fig. 2).
Figure 4: The sharp crossover between $\alpha=3/2$ to $\alpha=2$ scaling at $\omega\sim\sigma$, with $\sigma=10^{5}$ in this example. Here the blue line is $|\chi^{\prime\prime}|/\omega$ with $\chi^{\prime\prime}$ given by the exact expression, Eq. 15, The magenta dashed line is Eq.1 with $\alpha=3/2$, and the grey dashed line is $1/\omega^{2}$ with $\alpha=2$.
The limits $a\rightarrow 0$ ($u=0$) and $a\rightarrow\infty$ ($u$ finite) thus constitute two steady states where a frequency dependent susceptibility can be defined. In the former case the phonons are in equilibrium with the bath and have the Bose distribution $p^{0}=1/(e^{h\nu/kT}-1)$, while in the latter case the phonons have a non-equilibrium distribution.
Stretched exponential relaxation (Eq. 8) does not seem to have been tested experimentally for the phonon bottleneck in dilute spin systems, despite non-exponential relaxation being widely expected (and occasionally observed) following Ref. [12]. The many studies (see Ref. [16] for early work) have generally focused on temperature dependence, rather than time or frequency dependence, but there are at least two studies, by Standley and Wright [17], and by Roinel et al. [18], that tested Eq. 7 in the form of its solution for $t$:
$t=-\sigma\left(1-v\right)-(1+\sigma)\ln\left(v\right).$ (20)

### v3

@@ Abstract
The common observation of anomalous $1/f^{\alpha}$’ relaxation with $\alpha<2$ constitutes one of the enduring mysteries of condensed matter physics. Here it is shown that a $1/f^{\alpha}$ spectral density, with $\alpha=3/2$, can arise in the response of an ensemble of two–level systems coupled to a heat bath by means of a system of Bosonic quasiparticles. The model considered is the classic model of Faughnan and Strandberg of the phonon bottleneck, and the anomalous response is associated with an approximate non-equilibrium steady state of the phonons maintained by slow spin relaxation. The frequency dependence of the response to an applied field is calculated analytically, revealing the emergence, in the limit of a strong bottleneck, of $\alpha=3/2$ behaviour over a diverging range of frequencies. The application of this result to experimental systems is discussed and comparisons are drawn with other systems that exhibit anomalous relaxation.
I Introduction
@@
Exponential, or Debye, relaxation has a spectral density that tends to $1/f^{2}$ at high frequency $f$. However many real condensed matter systems – conducting [1, 2], dielectric [3, 4, 5] and magnetic [6, 7, 8, 9, 10, 11] – show power law behavior $\sim 1/f^{\alpha}$ with an exponent $\alpha$ that lies in the range $1\leq\alpha<2$. While a number of particular mechanisms for this behavior have been established (see e.g. Refs. [6, 7, 8, 9, 10] for the magnetic case), in general it remains a mystery as to why $1/f^{\alpha}$ behavior is so widespread. A common explanation [1] is that the power law is constructed from a wide range of exponential timescales, but in physical terms the range of timescales required (many decades) is usually unrealistic. An exception might be where the time scales are associated with an underlying broad range of length scales, as occurs in a critical or fractal system, but such instances cannot account for the majority of observations. This leads one to suspect that there are more fundamental causes at play, arising in non-equilibrium physics and quantum dynamics. In this context, simple and realistic models and experiments that produce well-defined exponents $\alpha$ are of evident interest to signpost the way to fundamental causes.
Straightforward mathematics, and more detailed studies of quantum systems [12], reveal two very general requirements for non–exponential relaxation. First, the system needs ‘memory’ of its relaxation, and second, memory is enhanced by disconnection of the system from its environment. In magnetism, at least partial disconnection from an environmental heat bath is always a factor to consider at sufficiently low temperature, because that connection is tyoically mediated by a rapidly decreasing population of increasingly long-lived phonons. This gives rise to the ‘phonon bottleneck’ which was characterised many years ago in dilute spin systems [13, 14, 15, 16, 17, 18, 19, 20, 21]. Unsurprisingly, the phonon bottleneck is associated with non-exponential relaxation [14, 16], but it is germane to ask the question, is it associated with any of the common analytic forms of non-exponential relaxation? Rather surprisingly, this question never seems to have been addressed in the literature.
The aim of this paper, therefore, is to answer the question stated above, with respect to the canonical model of the ‘phonon bottleneck’ due to Faughnan and Strandberg [14]. The question is simple, and an answer is immediately suggested by numerically solving the equations of Ref.[14] in the limit of a ‘strong’ bottleneck. The ensuing relaxation function, when digitally (‘Fast’) Fourier transformed, shows a clear $1/\omega^{3/2}$ dependence, as shown in Fig. 1 (here $\omega=2\pi f$)). The rest of the paper is concerned with understanding this result in terms of an analytic solution, exploring its relevance to real systems and discussing its implications for non-Debye relaxation in general.
Figure 1: Phonon bottleneck relaxation according to the model of ref. [14] with parameters $a=100,\sigma=500$ (parameters defined in Section II). Fast Fourier transform (FFT) of the numerically evaluated relaxation function $v(|t|)$ (blue dots) compared with a line (orange) of slope $3/2$ on a log-log scale, to illustrate an extensive range of $\sim 1/\omega^{3/2}$ relaxation.
In the following, the phonon bottleneck is reviewed (section II), the analytic result derived (section III), relation to real systems discussed (section IV) and conclusions drawn (section V).
II Review of the phonon bottleneck
It is well established that isolated magnetic moments (‘spins’) relax through three main processes: the direct process, the Raman process and the Orbach process [15, 17, 20]. The latter two processes tend to die away rapidly at low temperature leaving only the direct process, by which a phonon of appropriate energy is absorbed or emitted every time a spin flips. The spin-phonon interaction arises from a combination of lattice vibrations modulating the crystal electric field, and spin-orbit coupling in the case of electronic spins, or the nuclear quadrupole interaction in the case of nculear spins. In a perturbation treatment, matrix elements depend on phonon occupation numbers which become natural variables in the problem, alongside spin states (‘up’ or ‘down’ in the simplest case) [13]. If the magnetic heat capacity is much larger than the phonon heat capacity, which it typically is at low temperature, and if the phonon relaxation time starts to exceed the spin-lattice relaxation time, then magnetic relaxation drives phonon populations out of equilibrium, and magnetic relaxation is then determined by the rate of transfer of energy from the phonons to the bath. In the simplest case this manifests by a change in temperature dependence of the relaxation rate from $\sim T$ to $\sim T^{2}$ [15] (here $T$ is the bath temperature). The process can be treated approximately as a three-temperature problem (temperatures of spin system, phonons and bath) [17], but more generally the phonon and spin temperatures become poorly defined concepts, and it is more meaningful to work directly with phonon and spin state populations [13].
The canonical model of the phonon bottleneck due to Faughnan and Strandberg [14] makes two basic assumptions. The first is
$\frac{dN_{i}}{dt}=-\frac{dN_{j}}{dt}=-\rho A\left[N_{i}(p+1)-N_{j}p\right]$ (1)
where $N_{i}$ as the number of spins in the excited state, $N_{j}$ is the number of spins in the ground state, $p$ is the phonon occupation number in a narrow band at the frequency of the transition, $\rho$ is the number of phonon states, $A$ is a rate constant for stimulated emission or absorption, and the factor ‘1’ represents spontaneous emission. This equation may be manipulated into the form given in Ref. [14] using the additional condition that $N^{i}+N^{j}=N$, the number of spins. The second assumption is:
$\frac{d(p-p^{0})}{dt}=\frac{-1}{2\rho\Delta\nu}\frac{d(N_{i}-N_{j})}{dt}-\frac{(p-p^{0})}{\tau_{\rm ph}}.$ (2)
where (here and in what follows) the superscript 0 indicates equilibrium values, $\Delta\nu$ is the relevant band width and $T_{\rm ph}$ a phonon relaxation time that describes absorption by the heat bath or scattering into other phonon modes. This leads to two coupled equations of motion for the spin and phonon populations, which are conveniently rewritten in terms of dimensionless properties as follows.
@@ II Review of the phonon bottleneck
where
@@
and the variable $v$ is defined as
$v=\frac{(N_{i}^{0}-N_{j}^{0})-(N_{i}-N_{j})}{(N_{i}^{0}-N_{j}^{0})}.$ (6)
Here the parameter $a=\tau_{\rm sl}/\tau_{\rm ph}$ where $\tau_{\rm sl}$ is the spin-lattice relaxation time and the parameter $\sigma=b/a$. The latter can also be cast as:
@@
where the $c$’s are specific heats. A large value of $\sigma$ indicates that the energy exchange rate between spins and phonons greatly exceeds that between phonons and bath – hence it is a measure of the strength of the ‘phonon bottleneck’. Also it should be noted that the time $t$ here is measured in units of $\tau_{\rm sl}$.
The limiting situation of interest arises where $\sigma,a\gg 1$. This is the scenario illustrated in Fig. 1, which shows the numerical solution of the coupled equations for $a=100,\sigma=500$. Under such conditions the phonons are strongly bottlenecked but the spins relax relatively slowly, continually maintaining a far from equilibrium phonon population. This leads to an approximate steady state where the rate of change of the phonon population can be neglected in Eq. 2. Thus $|du/dt|\ll a$, and the right hand side of Eq. 4 may be set to zero. This reduces Eqs. 3, 4 to a single equation depending only on $\sigma$:
@@
This equation has an exact solution for time $t$:
$t=-\sigma\left(1-v\right)-(1+\sigma)\ln\left(v\right).$ (9)
Faughnan and Strandberg [14] showed that numerical solutions of both the coupled equations and the limiting equation Eq. 8 gave non-exponential relaxation, but the functional form of the relaxation was not explored. In addition they showed that Eq. 9 was a close approximation to the coupled equations for realistic parameters. In later work, Standley and Wright [16] found excellent agreement of these equations with experiment, as discussed further below. Later work on the phonon bottleneck largely developed the approach of Ref.[14] to include other factors, which include phonon diffusion [17], negative temperatures [18], and the more complex adaption of the model to describe the Orbach processes [19, 20].
III Analytic form of the relaxation
In order to establish notation and the meaning of relevant quantities, it is useful to briefly review some concepts of linear response theory. In linear response one typically imagines a system at equilibrium in an applied field $H$, such that the magnetization is $M^{0}=M(T\leq 0)=\chi_{T}H$. The field is then removed such that the system relaxes according to $M(t\geq 0)=R(t)$. Here $R(t)$ is a relaxation function, formally an even function of $t$. Clearly $R(0)=\chi_{T}$ and one can define a normalised function, $F(t)=R(T)/\chi_{T}$, whose Fourier transform is the spectral shape function $F(\omega)$. The response function $\chi(t)$ is defined by the convolution
$M(t)=\int_{-\infty}^{t}H(t^{\prime})\chi(t-t^{\prime})\,dt^{\prime}$ (10)
and it follows that $\chi(t<0)=0$ and $\chi(t\geq 0=-dR(t)/dt)$. The Fourier transform of $\chi(t)$ is the complex susceptibility $\chi(\omega)=\chi^{\prime}(\omega)+i\chi^{\prime\prime}(\omega)$, which connects the frequency dependent magnetization and field, $M(\omega)=\chi(\omega)H(\omega)$. It is possible to show that [22]
$\frac{2|\chi^{\prime\prime}|}{\chi_{T}\omega}=F(\omega).$ (11)
A time varying field of particular importance is $H(t)=he^{i\omega t}$. Then, if the reponse is $M(t)=me^{i\omega t}$, the ratio $\chi=m/h$ is equal to $\chi(\omega)$ derived from the Fourier transform of the response function. Explictly, the relationship is:
$\chi(\omega)=\int_{0}^{\infty}\chi(t)e^{-i\omega t}\,dt=\frac{m}{h}.$ (12)
The model of Faughnan and Strandberg (section II) may now be compared to these relationships. If $i$ denotes spin up, and $j$ denotes spin down, then the magnetic moment $\mathcal{M}=N_{i}-N_{j}$ (i.e opposite to the spin direction but this sign choice is ultimately not important) and the equilibrium magnetic moment is $\mathcal{M}^{0}=N_{i}^{0}-N_{j}^{0}$. It follows that $v=(\mathcal{M}^{0}-\mathcal{M})/\mathcal{M}^{0},$ a normalised measure of the deviation of magnetic moment, or magnetization, from its equilibrium value. For example, if a field is suddenly applied to an equlibrium unmagnetized system, then $v$ will evolve from $v=1$ to $v=0$ as the new equilibrium is established at $\mathcal{M}=\mathcal{M}^{0}$.
The alternative scenario of removing a field here leads to conceptual difficulties in physical terms. These are discussed in detail below, but for the time being one may proceed in a purely mathematical sense. Then, from the above definitions, it can be shown that $v(t\geq 0)=F(t)$, so the symmetrised $v(|t|)=F(t)$. In fact the limiting Eq. 9 can be inverted to give $v(t)$, and hence $F(t)$ exactly:
$v(|t|)=F(t)=-\frac{1}{1+\sigma}W\left(-\frac{\sigma\left(e^{-\sigma-|t|}\right)^{\frac{1}{1+\sigma}}}{1+\sigma}\right),$ (13)
where $W$ is the Lambert W-function.
The function Eq. 13 does not have a known Fourier transform but it can be integrated, giving
$F(\omega=0)=\sigma+2,$ (14)
a result that will be used below.
An analytic expression for $\chi(\omega)$ can now be derived. Thus Eq. 8 may be rewritten as a function of the relative magnetization $M/M^{0}$, such that
$1-\frac{M}{M^{0}}=\frac{1}{M^{0}}\frac{dM}{dt}\left(1+\sigma\frac{M}{M^{0}}\right),$ (15)
where $M^{0}=H\chi_{T}$. Now consider the case where $H=he^{i\omega t}$ and $M=me^{i\omega t}$. It follows that:
$\chi_{T}-\chi=i\omega\chi\left(1+\sigma\frac{\chi}{\chi_{T}},\right),$ (16)
from which
$\frac{\chi}{\chi_{T}}=\frac{i\left(-\sqrt{4i\sigma\omega-\omega^{2}+2i\omega+1}+iw+1\right)}{2\sigma\omega}$ (17)
consisting of the real and imaginary parts:
$\frac{\chi^{\prime}}{\chi_{T}}=\frac{-\omega+(1+(2+16\sigma(1+\sigma))\omega^{2}+\omega^{4})^{1/4}\sin\phi}{2\sigma w}$ (18)
$\frac{\chi^{\prime\prime}}{\chi_{T}}=\frac{1-(1+(2+16\sigma(1+\sigma))\omega^{2}+\omega^{4})^{1/4}\cos\phi}{2\sigma w}$ (19)
Figure 2: Frequency dependence of the derived complex susceptibility for $\sigma=500$. Upper plot: real part (orange dotted line, Eq. 18) and imaginary part (full blue line, Eq. 19) of the the analytic complex susceptibilty, showing three distinct frequency regimes. Lower plot: the corresponding function $2|\chi^{\prime\prime}|/\chi_{T}\omega$ in the region of the high frequency ‘crossover’, comparing the ‘exact’ Eq. 19 (blue line), the intermediate-frequency approximation Eq. 20 (orange dashed line) and the high frequency approximation $2/\omega^{2}$ (gray dashed line).
The real and imaginary parts of the susceptibility, on log-log scales, are illustrated in Fig. 2 where there is a very clear structure of three linear regimes. According to Eq. 19, for large for large $\omega<\sigma$,
$\frac{2|\chi^{\prime\prime}|}{\omega\chi_{T}}\rightarrow\frac{2}{\sqrt{\sigma}\,\omega^{3/2}},$ (20)
which picks out the ‘middle’ regime in Fig. 2, while for $\omega>\sigma$, high order terms in Eq. 19 ensure that the true asymptotic behavior is $1/\omega^{2}$, with a crossover to this form at $\omega\sim\sigma$. Thus $\alpha=3/2$’ scaling applies to the range $1/\sigma<\omega<\sigma$ and only the divergence of $\sigma$ would extend its range to infinity. Nevertheless, as illustrated in Fig. 2, for finite $\sigma$, the crossover is particularly sharp and has negligible effect on the power law scaling at smaller $\omega$.
The function $\chi(\omega)$ thus calculated gives the exact steady state response to an applied oscillating field. The function $2|\chi^{\prime\prime}|/\chi_{T}\omega$ unsurprisingly coincides with a numerical Fourier transform of the exact $F(t)$ at high frequency, but, somehwat unexpectedly, as $\omega\rightarrow 0$ it tends to the value $2\sigma+2$, in contrast to $\sigma+2$ given by Eq. 14. This perhaps suggests, for finite $\sigma$, there may not be a single response function that simultaneously covers both extreme cases of a step function applied field and a sinusoidal applied field.
Nevertheless, if $2|\chi^{\prime\prime}|/\chi_{T}\omega$ is adjusted by rescaling $\sigma\rightarrow\sigma/2$, such that it exactly coincides with the expected $F(\omega=0)$, then, as shown in Fig. 3 (which includes no fitted parameters), it provides an astonishingly close approximation to the numerical Fourier Transform of the exact $F(t)$ over the whole frequency range.
Figure 3: Phonon bottleneck relaxation in the limiting case with bottleneck parameter $\sigma=50$. Fast Fourier transform (FFT) of the exact relaxation function $v(|t|)$ (blue dots, Eq. 13) compared with the analytic $2|\chi^{\prime\prime}|/\chi_{T}\omega$ (magenta line, Eq. 19), showing distinct regimes of $\sim 1/\omega^{0},\sim 1/\omega^{3/2}$ and $2/\omega^{2}$ (Eq. 20) behaviour. In the analytic curve, the value of $\sigma$ is rescaled by a factor of 22 to exactly match $v(\omega=0)$. With this rescaling, and no fitting, the two curves are almost coincident at all frequencies, illustrating a rare instance where a $1/f^{\alpha}$’ spectral density (with $\alpha=3/2$) can be calculated analytically.
From this one may conclude that the true Fourier transform of $v(|t|)=F(t)$ also shows the emergence of a $1/\omega^{3/2}$’ regime, that will cover the whole domain of $\omega$ in the limit $\omega\rightarrow\infty$.
The three-slope function in frequency space may be anticipated (at least with the benefit of hindsight!) from Eq. 8. Thus, if $v$ can be approximated to either 1 (short times) or 0 (long times) in the denominator, then exponential relaxation is obtained, and this clearly translates through to the frequency response at large and small frequencies respectively. Also, the $1/\omega^{3/2}$ spectral function in frequency translates to $t^{-1/2}$ relaxation in the time domain and there is a hint of this in Eq. 8. Thus, at short times $v\rightarrow 1$, but for sufficiently large $\sigma$, there appears to be a region at $1-v\gg 1/\sigma$ where this equation approximates:
@@ III Analytic form of the relaxation
Despite these hints, the different regions of the relaxation curve are much less clearly distinguished in the time domain than they are in frequency space, and this likely explains why a $\sqrt{t}$’ dependence has not been noticed, or considered, before.
IV Application to real systems
In applying this result to real systems, three cases must be clearly distinguished.